Numerical Problem 9 n general, if Po is the prioe leval at the beginning of an n-year perlod and P is tha price level at the end of that period, the average annual rate of inflation over that perlod must satisfy the equation, The GDP dellatur in Econckand is 120 on Jarnuary 1, 2011. The dellalur rises to 241.4 by Jauary 1, 2016, and to 248.6 by January 1, 2017 What is the average annual rale of inllation over the 5 years between January 1, 2011, and January 1, 20167 In other words, what f inllation would lead to the price rise otserved over those 5 years? The average yearly rate of inflation that would lead to this price rise is 15 % Enter your snswer ees percentage rounded to one deams/place : What is the average annual rate of inflation over the 6 years between January 1, 2011, and January 1, 2017? In othar words, what constant ysary rate of Inflation would lead to the price rise observed over those 6 years? The average yearly ate of inflation that would lead to this price rise is 12.9 % Enter yo ra 5w r a pe ce rage ounded to one dec na place constant yearly rate the price tise observed Question is complete. Tap on the red indicalors to see incorrect answers
Answer #1

We are given the following equation

Pn/Po = (1 + x)^n where x is the inflation rate

1) 241.4 = 120(1 + x)^5

1 + x = (241.4/120)^0.2

x = 15.00%

Per year inflation rate on average is 15.0%.

2) 248.6 = 120(1 + x)^6

1 + x = (248.6/120)^0.167

x = 12.93%

Per year inflation rate on average is 12.9%.